OpenStudy (anonymous):

GCF ( factoring completely). 2a^2 - 2b^2

OpenStudy (anonymous):

\[2a^2-2b^2 = 2(a^2-b^2) = ?\] How do you think the difference inside the parentheses could be factored?

OpenStudy (anonymous):

2ab(a-b)????

OpenStudy (anonymous):

2(a-a) (b-b)????

OpenStudy (anonymous):

No, the last one would be equal to zero..

OpenStudy (anonymous):

So whats the answer!?

OpenStudy (anonymous):

Do you want the answer, or do you want to try to work it out step by step?

OpenStudy (anonymous):

Step by step

OpenStudy (anonymous):

Ok, let's play a game. If I told you that: \[a^2-b^2 \] can be written as a product of two parentheseses, the 1st of which is: \[(a+b)\] what do you think the other one would be?

OpenStudy (anonymous):

(A-b)

OpenStudy (anonymous):

Right! Bravo! So what would be the answer to you question?

OpenStudy (anonymous):

2(a-b)(a+b)

OpenStudy (anonymous):

Right! Well done!

OpenStudy (anonymous):

If its 6x^2-6^2 would it be. The same??

OpenStudy (anonymous):

6(x+y)(x-y)???? Or no

OpenStudy (anonymous):

You could say that: \[6x^2 - 6^2 = 6(x^2-6) = 6(x^2-\sqrt{6}^2)\] Then you can do the same thing we did before.

OpenStudy (anonymous):

But it's the gcf so can I say 3(x+y) 2(x-y)??

OpenStudy (anonymous):

??????

OpenStudy (anonymous):

I think it would be 6 in this case, but I'm not really sure